OpenStudy (anonymous):

I have no idea how to do this question can I get a little help? solve the equation for x. y=x-v/b

OpenStudy (anonymous):

do you know how to solve \[y=x-5\] for \(x\)? if the answer is "no" that is fine, i will explain it

OpenStudy (anonymous):

oh hold the phone it is actually \[y=\frac{x-v}{b}\]?

OpenStudy (anonymous):

yes that's what I meant I cant do that on my computer. but I don't know hav

OpenStudy (anonymous):

to solve it.

OpenStudy (anonymous):

ok lets solve \[\pi=\frac{x-5}{2}\]for \(x\) step 1) multiply by \(2\) to get rid of the fraction and get \[2\pi=x-5\] step 2) add \(5\) and get \(2\pi+5=x\) and now you are done because \(x\) is by itself

OpenStudy (anonymous):

now lets do the same thing with \[y=\frac{x-v}{b}\] step 1) multiply both sides by \(b\) and get \[by=x-v\] step 2) add \(v\) to both sides and get \[by+v=x\] and again we are done

OpenStudy (anonymous):

thanks I hate doing math.

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

well your better than me

OpenStudy (anonymous):

i have to part of my sentence